OpenStudy (anonymous):

Simplify the expression (2/(x-6))+4x

OpenStudy (anonymous):

you can take a GCF out.

OpenStudy (anonymous):

im sorry im still confused

OpenStudy (anonymous):

\[\huge \frac{ a }{ b }+\frac{ c }{ d } = \frac{ ad+bc }{ bd }\] Try using this.

OpenStudy (anonymous):

(4x^2-24x+2)/(x-6) im not sure how that helps, sorry.

OpenStudy (anonymous):

batman: \[\large\frac{ 2 }{ x-6 } + 4x = 2(\frac{ 1 }{ x-6 } + 2x)\]

OpenStudy (anonymous):

@eviI but what about the x-6

OpenStudy (anonymous):

FOUND IT FINALLY, thanks for all your help guys. sorry i can be a knuckle head sometimes.

OpenStudy (anonymous):

ye